# CaveAgent: Transforming LLMs into Stateful Runtime Operators

### v1

@@ 3.1.1 Variable and Function Injection
To bridge the gap between large language models and executable environments, we introduce a unified abstraction for Variable and Function injection. CaveAgent treats Python objects and functions as first-class citizens within the runtime environment to ensure object-oriented interactions. This mechanism consists of two key components: metadata extraction for the semantic stream’s context and direct object injection into the runtime namespace.
Descriptive Abstraction
Each injectable entity is wrapped in a container that automatically extracts its metadata. For functions, this includes the signature, type hints, and docstrings; for variables, it includes the name, type, and an optional description. Formally, a function $f$ is represented as a tuple $(n_{f},\sigma_{f},d_{f})$, where $n_{f}$ is the function name, $\sigma_{f}$ is the signature derived from inspection, and $d_{f}$ is the documentation. Similarly, a variable $v$ is represented as $(n_{v},\tau_{v},d_{v})$, where $\tau_{v}$ is the type. This metadata is aggregated and injected into the system prompt, providing the model with a clear "API reference” of available capabilities without exposing implementation details or raw values. For example, an injected data processing object might be presented to the model as:
name: processor
type: DataProcessor
description: A utility for sorting and filtering data collections.
methods: process(data: list) -> list
Namespace Injection
Critically, injection goes beyond mere description. Upon initialization, the runtime maps these entities directly into the namespace of the underlying execution engine (e.g., IPython). This means that if a function add or an object processor is injected, they become immediately available as global symbols in the execution environment. This design enables Object-Oriented Interaction and Stateful management. Instead of stateless function calls (e.g., tool: "sort", args: {data: ...}), the model can invoke methods on stateful objects directly (e.g., processor.process(data)). This significantly enhances composability, as the model can chain method calls and manipulate object attributes naturally, mirroring standard programming practices rather than rigid API request-response cycles.
After variable and function injection, CaveAgent interacts with the environment via executable Python programs leveraging native Python syntax for robust parsing and utilizing control flow (loops, conditionals) with stateful data passing to handle multi-step logic. Unlike text-based paradigms, CaveAgent allows for lossless manipulation of complex data structures throughout the interaction. Consequently, the agent delivers the final output not as a textual approximation, but as a valid, native Python object guaranteed to match the expected type, enabling seamless integration with downstream applications.
@@ 3.1.2 Dynamic Context Synchronization
While the dual-stream architecture decouples reasoning from state storage, effective collaboration requires a regulated information flow between the Semantic Stream and the Runtime Stream. We implement a dynamic synchronization mechanism to ensure the agent remains aware of the runtime state without overwhelming its context window.
In our framework, the Semantic Stream is "blind" to the Runtime Stream by default. Visibility is achieved explicitly via execution outputs. To inspect the state $\mathcal{S}_{t}$ (e.g., the content of a variable), the agent must generate code to print a summary (e.g., print(df.head())). This design enforces an Active Attention Mechanism: the agent consciously selects which part of the massive runtime state is relevant to the current reasoning step, pulling only that slice from the runtime, the external memory, into the token context.
To prevent "Context Explosion" caused by accidental verbose outputs (e.g., printing a million-row list), we introduce an Observation Shaping layer. The runtime captures standard output and subjects it to a length constraint function $\tau(\cdot)$.
$o_{t}=\tau(s_{t})=\begin{cases}\text{stdout}_{t}&\text{if }|\text{stdout}_{t}|\leq L_{\max}\\ \text{Error}_{\text{size}}(\text{len})&\text{if }|\text{stdout}_{t}|>L_{\max}\end{cases}$ (4)
When the output exceeds $L_{\max}$, instead of truncating silently, the system injects a specific meta-instruction prompting the agent to revise its code (e.g., "Output exceeded limit, please use summary methods"). This feedback loop teaches the agent to interact with the persistent state efficiently, favoring information which is concise and most-relevant, over verbose raw data dumps.
@@ 3.1.3 Security Check via Static Analysis
CaveAgent mitigates code execution risks via Abstract Syntax Tree (AST)-based static analysis, enforcing security policies without compromising flexibility. We parse code $c_{t}$ into a tree $\mathcal{T}_{c_{t}}$ and validate it against a policy $\Pi=\{r_{1},\dots,r_{k}\}$, where $\text{Valid}(c_{t})\iff\forall r\in\Pi,r(\mathcal{T}_{c_{t}})=\emptyset$. The modular rule set includes (example):
  • ImportRule: Blocks unauthorized modules (e.g., os, subprocess).
  • FunctionRule: Prohibits dangerous calls (e.g., eval(), exec()).
  • AttributeRule: Prevents sandbox bypass via internals (e.g., __builtins__).
Structured Error Feedback. Violations trigger structured observations rather than system crashes. For instance, a SecurityError is returned to the semantic stream, enabling the agent to self-correct (e.g., replacing eval() with safe tools) and ensuring interaction continuity.
@@ 4 Experiments
In this section, validate CaveAgent by answering four questions:
@@
To verify the CaveAgent’s basic function-calling capabilities on standard function-calling tasks, we employ two widely-adopted benchmarks in Agentic tool use: Tau2-bench (yao2024tau) and the Berkeley Function Calling Leaderboard (BFCL) (bfcl).
Models.
We evaluate a wide spectrum of State-of-the-Art (SOTA) LLMs to benchmark our performance, ensuring a comprehensive coverage of different architectures (e.g., dense vs. MoE) and model scales. The model suite includes:
  • DeepSeek-V3.2: The latest iteration of the DeepSeek series, featuring a Mixture-of-Experts (MoE) architecture with 685B parameters (37B active). Setting: Temperature set to $0.2$ to ensure stable code generation.
  • Qwen3 Coder 30B: A specialized code-centric model built on the Qwen3 architecture. It utilizes a highly efficient MoE design with 30B parameters (3B active). Setting: Configured with a temperature of $0.2$ for stable output generation.
  • Kimi K2 0905: A large-scale MoE model with 1000B parameters (32B active), designed for long-context interactions. Setting: We adopt the official recommended temperature of $0.6$.
  • Claude Sonnet 4.5: The SOTA model of Claude-series from Anthropic. Setting: Temperature is set to $0.2$ for stable code generation.
  • GPT-5.1: An evolution of the GPT-series. Setting: We utilize the default temperature of $1.0$, as this is the only supported value for the current snapshot.
  • Gemini 3 Pro: Known for its massive context window and native multimodal reasoning. Setting: Configured with "Low thinking” reasoning mode and a temperature of $1.0$, adhering to official recommendations.
For each backbone model, we conduct a comparative analysis between its native function-calling mechanism and our proposed CaveAgent framework. Crucially, within the CaveAgent workflow, the LLM is repurposed solely as a text generation engine, referred to as the semantic stream in our framework, bypassing its internal function-calling modules. We run each model using the standard API offered by the model provider.
@@ 4.1.1 Results on Tau2-bench
Tau2-bench (yao2024tau) is a comprehensive benchmark designed to evaluate the dynamic tool-use capabilities of LLM-based agents in realistic, multi-turn conversational scenarios. Unlike static evaluation sets that focus on single-turn intent detection, Tau2-bench necessitates that the agent interacts with a simulated user to achieve complex goals (e.g., modifying a flight reservation or processing a retail refund) while maintaining consistency across multiple turns. Following the original Tau2-bench paper, we focus on two primary domains within the benchmark: Airline and Retail. These domains are challenging since they require the agent to accurately track user constraints, database states, and policy regulations throughout the dialogue history.
Experimental Setup.
To ensure a rigorous comparison, we strictly follow the evaluation protocols of Tau2-bench. Specifically, we utilize DeepSeek V3 as the user simulator for all experiments to generate diverse and coherent user responses. To mitigate variance in generation, each model is tested three times for each domain. The reported results represent the respective scores and average scores of these three independent runs.
Evaluation of CaveAgent.
Since CaveAgent executes Python code rather than JSON tool calls, we employ runtime instrumentation to capture function invocations. Wrapper functions intercept each function call, recording function names and arguments before delegating to the underlying implementation. The captured invocation sequence is compared against ground-truth actions using identical evaluation criteria applied to JSON-based agents, ensuring a fair cross-paradigm comparison based on which functions were called with what arguments.
@@
The quantitative results on Tau2-bench are summarized in Table 1. The key insights include: (1). CaveAgent consistently outperforms the standard JSON-based function calling paradigm across 11 out of 12 experimental settings, covering both open-source and proprietary models ranging from 30B to over 1000B parameters. Significant improvements are observed in most SOTA models like DeepSeek-V3.2 and Gemini 3 Pro (averaging +5.3% and +6.1% respectively), demonstrating that our framework breaks the performance ceiling of even the most capable semantic reasoners by offloading state management to a deterministic and error-free code runtime. (2). CaveAgent shows superiority in state-intensive scenarios. We observed that the performance advantage is markedly amplified in the Retail domain compared to Airline. Retail tasks in Tau2-bench typically involve complex transaction modifications and policy checks, which require maintaining high-fidelity state consistency across turns. The standard paradigm suffers from serialization overhead here, leading to hallucinations. In contrast, CaveAgent achieves double-digit gains in Retail for models like Qwen3 and Kimi K2. This validates our hypothesis that Stateful Runtime Management effectively eliminates errors caused by the repetitive text-based serialization of complex data objects (e.g., shopping carts or refund policies). We provided a detailed agent trajectory analysis about the reason of CaveAgent’s outstanding performance in Retail tasks in Appendix F.1. (3). CaveAgent unlocks the potential in code-centric models. Most notably, the smaller, code-specialized Qwen3-Coder (30B) exhibits the largest relative improvement (+13.5% in Retail), enabling it to rival the performance of significantly larger generic models. This confirms that CaveAgent effectively leverages the inherent coding proficiency of LLMs. By decoupling the semantic stream from the runtime stream, our approach allows code-centric models to focus on logic generation rather than struggle with verbose context tracking, thereby maximizing the utility of limited parameters.
@@ 4.1.2 Results on BFCL
While Tau2-bench evaluates the capability of maintaining long-term state, it is equally critical to assess the agent’s precision in atomic function executions. To this end, we employ the Berkeley Function Calling Leaderboard (BFCL) (bfcl), a widely recognized benchmark for quantifying the accuracy of LLM tool invocation.
Benchmark Overview.
BFCL constructs a rigorous evaluation environment consisting of approximately 2,000 question-function-answer pairs derived from real-world use cases. The dataset is designed to test models across varying levels of complexity. Key evaluation categories include:
  • Simple Function: Represents the fundamental evaluation scenario where the model is presented with a single function definition and must generate a unique invocation with correct arguments and results.
  • Multiple Function: Assesses the model’s selection capability. The model is provided with a candidate set of 2 to 4 function definitions and must identify and execute the single most appropriate function that addresses the user’s query, filtering out irrelevant tools.
  • Parallel Function: Evaluates the ability to execute concurrent/parallel actions within a single turn. The model must decompose a complex user query (spanning one or multiple sentences) into multiple distinct function calls, invoking them simultaneously to optimize efficiency.
  • Parallel Multiple Function: The most challenging category, combining tool selection with parallel execution. The model is confronted with a larger pool of function definitions and must determine both the correct subset of tools to use and the frequency of their invocation (zero or more times) to fully resolve the request.
Notably, we use Executable Evaluation (Functional Correctness), rather than Abstract Syntax Tree (AST) evaluation since AST is not directly applicable in our CaveAgent framework due to the lack of explicit JSON schema. We execute the generated code in a controlled environment and compares the execution result against the ground truth. This validation ensures that the function call triggers the correct behavior in real-world applications.
Table 2: Detailed Performance comparison on BFCL Benchmark (3 Runs). Data is presented in score/total format. The Avg. columns indicate the average overall percentage across 3 runs. Bold indicates CaveAgent outperforms Function Calling. Simp. means Simple Function, Mult. means Multiple Function, Para. means Parallel Function, P-M. means Parallel Multiple Function, Ov. means the overall performance across four function categories, and Avg. means the average performance across three runs.
Model Run Function Calling CaveAgent
Simp. Mult. Para. P-M. Ov. Avg.(%) Simp. Mult. Para. P-M. Ov. Avg.(%)
Open Source
Closed Source
R3 387/400 190/200 184/200 184/200 945/1000 94.4 385/400 189/200 184/200 186/200 944/1000 94.4 ( 0.0)
R3 384/400 190/200 188/200 182/200 944/1000 94.3 380/400 194/200 184/200 185/200 943/1000 94.3 ( 0.0)
@@
The results in Table 2 highlight the atomic precision of CaveAgent in single-turn scenarios. Key observations include: (1). The results for DeepSeek-V3.2 (w/o prompt) reveal a critical insight. We hypothesize that due to its training emphasis on reasoning regarding tool dependencies (liu2025deepseek), DeepSeek-V3.2 exhibits a strong inductive bias toward sequential execution, causing it to fail in parallel-calling scenarios under the standard JSON paradigm (53.1% accuracy). To ensure a fair comparison, we added explicit prompting to the system prompt of DeepSeek V3.2 to "force” parallel execution. In stark contrast, CaveAgent achieves SOTA performance (94.0%) without any prompt intervention. This demonstrates a unique advantage of our paradigm: by utilizing Python code, CaveAgent naturally supports parallel execution (e.g., via independent lines of code) while simultaneously preserving the capacity to reason about inter-tool dependencies, resolving the conflict between reasoning depth and execution parallelism that standard JSON approaches struggle. (2). The 30B parameter Qwen3-Coder, when equipped with our framework, achieves a 94.4% average score, outperforming the much larger proprietary GPT-5.1 (89.6%) and matching Claude Sonnet 4.5. We attribute this to that CaveAgent unlocks more potential of small LLMs via effectively leveraging the inherent coding proficiency of LLMs.
For most SOTA models like Claude Sonnet 4.5 and Gemini 3 Pro, CaveAgent performs on par with the standard baseline (94.3%), with negligible variance. We attribute this plateau to benchmark saturation. Current SOTA models have likely reached the upper limit of the BFCL dataset, where remaining errors stem from ambiguous natural language queries or ground-truth noise rather than model incapacity. Since BFCL focuses strictly on single-turn intent detection without the complexity of state maintenance, the "ceiling" is hit relatively quickly. It is important to emphasize that Tau2-bench and BFCL serve primarily to validate the basic function-calling capabilities of CaveAgent. However, the true superiority of our proposed paradigm lies in tasks demanding more advanced tool-use—specifically, the manipulation of complex data objects over long-horizon tasks. Consequently, existing benchmarks are insufficient to fully capture the stateful management capability of CaveAgent. In the following section, we utilize our hand-crafted cases to provide a deeper and more rigorous assessment of CaveAgent’s capabilities in long-horizon stateful management.
@@
To evaluate CaveAgent’s stateful runtime management capabilities, we design a benchmark targeting multiple complementary dimensions of state manipulation that existing function-calling benchmarks fail to address. The benchmark tests an agent’s ability to read, modify, and persist variables across multiple conversational turns. We divide the measurement of stateful management into three categories: Python type proficiency, capability of multi-variable manipulation, and robustness in multi-turn, long-horizon interaction. A unifying design principle is programmatic validation: rather than parsing text outputs or relying on heuristic matching, we directly inspect runtime state after execution, verifying exact values, object attributes, and data structure contents against ground-truth expectations. This enables precise, unambiguous evaluation and demonstrates a key advantage of CaveAgent’s architecture: agent behavior becomes programmatically verifiable, opening pathways for automated evaluation and reinforcement learning with accurate reward signals.
For each dimension, we manually curate multiple test cases, each consisting of multiple natural language queries and an initial variable state, where the queries are linearly dependent. The agent sequentially manipulates the variable according to the queries, after which we retrieve the resulting variable for validation. A query is considered successful if the value of the output variable aligns with our expectations (see Appendix D for the details about the test cases). To isolate core state management capabilities, we craft queries with unambiguous requirements and explicit expected outcomes, ensuring that failures reflect genuine limitations in state tracking rather than instruction misinterpretation. Multiple queries per case further measure long-horizon state persistence and numerical precision across multi-step operations. For each dimension, we select four models to conduct this experiment: QWen3 Coder, Kimi K2 0905, Deepseek V3.2 and Gemini3 Pro. We report the success rate computed by the number of successful queries/total number of queries, as shown in Table 3.
@@
Kimi K2 0905 Preview 100 100 100 100 100 100 100 100 100 100 90.0 100 95.0
Gemini 3 Pro Preview 100 100 100 100 100 100 100 100 100 100 97.5 100 98.7
4.2.1 Type Proficiency
Type Proficiency aims to evaluate an agent’s ability to manipulate variables across a spectrum of Python types:
  • Simple Types: Fundamental operations on Python primitive types including integers, floating-point numbers, strings, booleans, lists, and dictionaries.
  • Object Types: Interaction with user-defined class instances, including attribute access and modification, method invocation, and state tracking across object lifecycles.
  • Scientific Types: Proficiency with data science primitives commonly used in computational workflows: pandas DataFrames, pandas Series, and NumPy ndarrays. Operations include column creation, filtering, sorting, element-wise transformations, aggregations, and cross-type interactions (e.g., storing array computations as DataFrame columns).
The results yield uniformly high scores (96.5%–100%), validating that code-based manipulation of complex types—including DataFrames, ndarrays, and custom objects—is tractable for current LLMs.
4.2.2 Multi-Variable
The Multi-Variable benchmark evaluates how state management accuracy changes when number of variables scale up.
The benchmark comprises five tiers with 5, 10, 15, 20, and 25 concurrent variables, systematically testing the agent’s working memory capacity and ability to perform coordinated state manipulation. Each tier contains 5 multi-turn conversations, and each conversation contains 3 turns, yielding 15 evaluation points per variable tier. The results show no systematic degradation as variable count scales to 25, with top models maintaining 100% accuracy throughout, demonstrating that concurrent state management scales effectively within CaveAgent’s architecture.
4.2.3 Multi-Turn
The Multi-Turn benchmark assesses an agent’s ability to read, modify, and persist variables across extended interactions—a critical capability for real-world deployments requiring the tracking of cumulative state changes. The benchmark comprises two domain-specific scenarios, each spanning 40 turns across two conversations:
  • Smart Home: Simulates a home automation environment where the agent manages devices (e.g., lighting, thermostats) via natural language. This scenario tests the agent’s ability to interpret intent and maintain device state consistency as commands accumulate.
  • Financial Account: Simulates banking operations such as transfers and inquiries. This scenario specifically targets numerical precision—ensuring calculation accuracy over multi-step operations—and stateful reasoning within a growing transaction history.
Collectively, these scenarios evaluate long-horizon state persistence, measuring whether the agent can reliably modify and track program state without drift as the conversation length increases. The results reveal the most meaningful differentiation between models. Long-horizon state persistence across 40 turns proves challenging: while DeepSeek-V3.2 maintains perfect accuracy, other models exhibit degradation, particularly on Smart Home scenarios requiring object state consistency. This suggests that accumulated state tracking over extended interactions remains the frontier capability for stateful agents. The consistently high accuracy across top models validates our central thesis: when LLMs interact through code with persistent runtime state, reliable and verifiable agent behavior becomes achievable. Notably, we restrict this evaluation to CaveAgent, as the fine-grained programmatic verification of Python objects is fundamentally incompatible with the text-based outputs of the JSON function-calling paradigm. Nevertheless, the near-perfect performance exhibited by CaveAgent independently substantiates the robustness of our stateful management.
@@ 4.2.3 Multi-Turn
As a complementary experiment, we evaluate CaveAgent’s advantages in context engineering and token efficiency against traditional JSON-based function calling. We benchmark CaveAgent across three domains: IoT device control, financial portfolio analysis, and e-commerce operations. This benchmark specifically targets scenarios requiring logically interdependent tool operations, creating "check → decide → act" cycles where multiple tool calls depend on prior results. We evaluated both methods using DeepSeek V3.2, measuring success rate, execution steps, and token consumption to isolate and quantify the efficiency gains attributable to the architectural shift from iterative JSON dispatching to native code generation.
Figure 5 shows the full results of this study, and Table 4 summarizes the performance improvement by comparing the summed performance metrics across three domains. The results demonstrate that CaveAgent achieves 28.4% lower total token consumption (504K vs. 704K) while improving task success rate from 94.6% to 100%. The efficiency gain stems from reducing interaction turns. Traditional function calling requires separate request-response cycles for each dependent operation, causing prompt tokens to accumulate as conversation history grows with each turn. CaveAgent instead generates Python code that resolves multiple dependencies in a single execution, reducing total steps from 236 to 145 and consequently cutting prompt tokens by 32.7%. More importantly, the Stateful Management property of CaveAgent naturally reduces the token overhead in multi-turn interaction. This is because CaveAgent manipulates persistent objects via variable references rather than repeatedly serializing full data states into text, as required by stateless process-oriented paradigms.
Notably, CaveAgent consumes 36.3% more completion tokens, since Python code with loops and conditionals is more verbose than JSON schemas. However, prompt tokens dominate overall consumption and accumulate across turns, while completion tokens only account for very small proportion of total token consumption.
@@
To assess the practical benefits of stateful runtime management, we evaluate three agent architectures on a data-intensive benchmark comprising 30 tasks across data query, statistical analysis, and visualization. The benchmark uses stock market data from Apple and Google (2020–2025, Yahoo Finance, retrieved from https://finance.yahoo.com). To ensure consistency, all three task categories were equipped with identical data retrieval functions. CodeAct Style replicates standard code-execution agent behavior by disabling CaveAgent’s variable injection and retrieval. JSON-based Function Calling operates without code execution, relying solely on tool outputs fed back to the model. The results are shown in Table 5.
Data Query Tasks
CaveAgent achieved 100% accuracy while consuming only 123K tokens by storing query results directly in runtime variables, thereby bypassing prompt context accumulation. In contrast, CodeAct Style (80%, 250K tokens) and Function Calling (80%, 295K tokens) both required serializing complete datasets into the conversation history, either through printed output or tool responses, resulting in context overflow failures on high-volume queries. Notably, both baseline methods failed on identical tasks involving large result sets that exceeded the model’s context limit.
Data Analysis Tasks
For tasks requiring statistical computation (e.g., volatility, correlation, Sharpe ratio), CaveAgent and CodeAct Style both achieved 100% accuracy with comparable token consumption ( 116–119K), demonstrating that code execution is essential for programmatic analysis. However, Function Calling achieved only 10% accuracy while consuming over 1.3M tokens. Without code execution capabilities, the model could only succeed on a trivial counting task; all computational tasks failed as the model attempted to infer statistics from raw data rather than compute them programmatically.
Visualization Tasks
These tasks required generating ECharts configurations containing both chart specifications and underlying data arrays. CaveAgent maintained 90% accuracy (405K tokens) by retrieving computed chart data from runtime variables without context serialization. CodeAct Style achieved 40% accuracy but consumed approximately 1M tokens, as generated visualizations must be printed to the conversation for output extraction. Function Calling achieved only 30% accuracy (662K tokens)—lower than CodeAct despite fewer total tokens—because earlier context overflow caused complete failures before task completion.
Discussion
These findings demonstrate that stateful runtime management provides substantial efficiency gains for data-intensive agent tasks. By decoupling intermediate computational state from the prompt context, CaveAgent avoids the token accumulation that causes context overflow in conventional architectures. This advantage becomes increasingly crucial as task complexity and data volume scale, suggesting that persistent runtime environments represent a promising direction for building robust agentic systems capable of handling real-world data processing workloads.

### v2

@@ 2.1.1 Variable and Function Injection
CaveAgent treats Python objects and functions as first-class citizens within the runtime. Each injectable entity is wrapped in a container that automatically extracts metadata: signatures, type hints, and docstrings for functions; names, types, and descriptions for variables. This metadata is aggregated into the system prompt as a lightweight “API reference,” while the actual objects are mapped directly into the execution engine’s namespace as global symbols. This design enables Object-Oriented Interaction: instead of stateless JSON calls (e.g., tool: "sort", args: {...}), the model invokes methods on stateful objects directly (e.g., processor.process(data)), chaining method calls and manipulating attributes naturally. The agent interacts via executable Python programs with native control flow (loops, conditionals) and stateful data passing, delivering final output as a native Python object rather than a textual approximation.
@@ 2.1.2 Dynamic Context Synchronization
The Semantic Stream is “blind” to the Runtime Stream by default. To inspect runtime state, the agent must explicitly generate code (e.g., print(df.head())), enforcing an Active Attention mechanism that selectively pulls only relevant slices into the token context. To prevent context explosion from verbose outputs, an Observation Shaping layer enforces a length constraint: when output exceeds $L_{\max}$, the system returns a structured error prompting the agent to use summary methods. This feedback loop guides efficient interaction with persistent state.
@@ 2.1.3 Security Check via Static Analysis
CaveAgent mitigates code execution risks via AST-based static analysis with modular policy rules: ImportRule (blocking unauthorized modules), FunctionRule (prohibiting dangerous calls like eval()), and AttributeRule (preventing sandbox bypass). Violations return structured errors to the semantic stream, enabling self-correction without breaking interaction continuity.
2.1.4 Runtime-Integrated Skill Management
Figure 4: CaveAgent’s Skill Management extends the Agent Skills open standard with runtime injection. Standard skills provide text prompts; CaveAgent skills additionally export functions, variables, and types into the persistent runtime.
CaveAgent extends the Agent Skills open standard [^2] https://agentskills.io, a portable specification adopted across the AI ecosystem including Claude Code, Gemini CLI, and Cursor. by introducing an injection.py module alongside the standard SKILL.md file. While standard skills provide text-based prompts that guide LLM behavior, CaveAgent skills additionally export Functions, Variables, and Type definitions that are injected directly into the persistent runtime upon activation. The framework employs progressive disclosure: skill metadata (name and description) is loaded at startup for routing decisions, while full instructions and runtime injections are loaded on-demand when the agent invokes activate_skill(). This bridges declarative skill definitions with CaveAgent’s stateful runtime paradigm: skills become executable capabilities, not merely textual instructions. As shown in Figure 4, domain expertise is packaged as both human-readable instructions for the language model and machine-executable artifacts for the runtime.
Table 1: Comparison of Standard Agent Skills vs. CaveAgent Skills.
Property Standard Skills CaveAgent Skills
Skill format SKILL.md SKILL.md + injection.py
LLM interaction Reads text prompts Operates on injected objects
Capability delivery Textual instructions Executable runtime artifacts
Data flow Text-in, text-out Object-in, object-out
@@ 3 Experiments
In this section, we validate CaveAgent by answering four questions:
@@
We evaluate on Tau2-bench (yao2024tau) and BFCL (bfcl) using six SOTA LLMs: DeepSeek-V3.2 (685B MoE), Qwen3 Coder (30B MoE), Kimi K2 0905 (1000B MoE), Claude Sonnet 4.5, GPT-5.1, and Gemini 3 Pro. For each model, we compare its native function-calling mechanism against CaveAgent, where the LLM serves solely as a text generation engine bypassing internal function-calling modules. Temperature settings follow official recommendations.
@@ 3.1.1 Results on Tau2-bench
Tau2-bench evaluates multi-turn tool use in realistic conversational scenarios (Airline and Retail domains), requiring agents to maintain consistency across turns. We follow the original evaluation protocols with DeepSeek V3 as user simulator, testing each model three times per domain. Since CaveAgent executes Python code, we employ runtime instrumentation with wrapper functions to capture and compare function invocations against ground truth, ensuring fair cross-paradigm evaluation.
@@
The results on Tau2-bench are summarized in Table 2. Key findings include:
(1). CaveAgent outperforms JSON-based function calling in 11 out of 12 settings across models from 30B to over 1000B parameters, with consistent improvements for DeepSeek-V3.2 (+5.3%) and Gemini 3 Pro (+6.1%), showing that offloading state management to a deterministic code runtime improves performance.
(2). Gains are amplified in state-intensive Retail scenarios, where complex transaction modifications require maintaining state consistency across turns. CaveAgent achieves double-digit gains for Qwen3 and Kimi K2, validating that Stateful Runtime Management reduces serialization-induced errors (detailed trajectory analysis in Appendix G.1).
(3). The code-specialized Qwen3-Coder (30B) exhibits the largest improvement (+13.5% in Retail), rivaling larger models. CaveAgent leverages the inherent coding proficiency of LLMs, allowing code-centric models to focus on logic generation rather than verbose context tracking.
@@ 3.1.2 Results on BFCL
To complement Tau2-bench’s multi-turn evaluation, we assess atomic function-calling precision on the Berkeley Function Calling Leaderboard (BFCL) (bfcl), which comprises $\sim$2,000 question-function-answer pairs across four categories of increasing complexity: simple, multiple, parallel, and parallel-multiple function calling. We use Executable Evaluation (functional correctness) by executing generated code and comparing results against ground truth. The summary is shown in Table 3 (detailed per-run results in Appendix Table 7).
Table 3: BFCL Benchmark Summary (avg. of 3 runs). Bold indicates CaveAgent outperforms Function Calling.
Model FC Avg.(%) CaveAgent Avg.(%) $\Delta$
DeepSeek-V3.2 (685B) 86.9 94.0 +7.1
DeepSeek-V3.2 (w/o prompt) 53.1 94.0 +40.9
Qwen3-Coder (30B) 89.8 94.4 +4.6
Kimi-K2-0905 (1000B) 89.2 94.7 +5.5
Claude Sonnet 4.5 94.4 94.4 0.0
GPT-5.1 89.6 88.9 -0.7
Gemini 3 Pro 94.3 94.3 0.0
@@
DeepSeek-V3.2 without explicit parallel-execution prompting achieves only 53.1% under the JSON paradigm due to its strong inductive bias toward sequential execution (liu2025deepseek). CaveAgent achieves 94.0% without any prompt intervention, as Python code naturally supports parallel execution via independent statements while preserving inter-tool dependency reasoning. The 30B Qwen3-Coder with CaveAgent (94.4%) outperforms the much larger GPT-5.1 (89.6%) and matches Claude Sonnet 4.5, demonstrating that CaveAgent leverages the coding proficiency of smaller LLMs. For SOTA models already at near-ceiling performance (Claude Sonnet 4.5, Gemini 3 Pro), gains are negligible since remaining errors stem from ambiguous queries rather than model incapacity. The advantages of our paradigm are most apparent in tasks requiring manipulation of complex data objects over long-horizon interactions, which we assess next.
@@
We design a benchmark targeting dimensions of state manipulation that existing benchmarks do not address, measuring an agent’s ability to read, modify, and persist variables across turns. A key design principle is programmatic validation: we directly inspect runtime state after execution against ground-truth expectations, rather than parsing text outputs. For each dimension, we curate test cases with linearly dependent queries and initial variable states (see Appendix D). Results are shown in Table 4.
@@
Kimi K2 0905 100 100 100 100 100 100 100 100 100 100 90.0 100 95.0
Gemini 3 Pro 100 100 100 100 100 100 100 100 100 100 97.5 100 98.7
Type Proficiency
evaluates manipulation of Python primitives, user-defined class instances, and scientific types (DataFrames, ndarrays). Results yield uniformly high scores (96.5%–100%), validating that code-based manipulation of complex types is tractable for current LLMs.
Multi-Variable
tests how accuracy scales with 5–25 concurrent variables across five tiers (15 evaluation points each). Top models maintain 100% accuracy throughout, demonstrating that concurrent state management scales effectively within CaveAgent’s architecture.
Multi-Turn
assesses state persistence across 40-turn interactions in two scenarios: Smart Home (device state consistency) and Financial Account (numerical precision over multi-step operations). While DeepSeek-V3.2 maintains perfect accuracy, other models exhibit degradation on long-horizon state tracking. The consistently high accuracy across top models validates our thesis: when LLMs interact through code with persistent runtime state, reliable and verifiable agent behavior becomes achievable.
@@
We evaluate token efficiency across three domains (IoT, finance, e-commerce) with logically interdependent tool operations using DeepSeek V3.2. Figure 5 and Table 5 show results. CaveAgent achieves 28.4% lower total token consumption (504K vs. 704K) while improving success rate from 94.6% to 100%. The gain stems from resolving multiple dependencies in single code executions, reducing steps from 236 to 145 and prompt tokens by 32.7%. While CaveAgent consumes 36.3% more completion tokens (code is more verbose than JSON), prompt tokens dominate overall consumption and accumulate across turns.
@@
We evaluate three architectures on a data-intensive benchmark comprising 30 tasks across data query, analysis, and visualization using stock market data (Apple and Google, 2020–2025). CodeAct Style disables CaveAgent’s variable injection/retrieval; JSON-based Function Calling operates without code execution. Results are shown in Table 6.
Results.
On Data Query, CaveAgent achieved 100% accuracy (123K tokens) by storing results in runtime variables, while both baselines failed at 80% due to context overflow from serializing large datasets. On Data Analysis, CaveAgent and CodeAct both achieved 100% with comparable tokens ($\sim$116–119K), but Function Calling managed only 10% (1.3M tokens) without code execution. On Visualization, CaveAgent achieved 90% (405K tokens) by retrieving chart data from runtime variables; CodeAct reached 40% (1M tokens) and Function Calling 30% (662K tokens). These results demonstrate that decoupling intermediate state from prompt context avoids the token accumulation causing context overflow in conventional architectures, with advantages growing with task complexity and data volume.